a father is three times as old as his son. after twelve years, his age will be twice as that of his son then. find their present ages.
Answer by stanbon(75887)

a father is three times as old as his son. after twelve years, his age will be twice as that of his son then. find their present ages.
----------
Equations:
f = 3s
f+12 = 2(s+12)
----
Substitute for "f" and solve for "s":
3s+12 = 2s+24
s = 12 yrs (the son's age)
----
Solve for "f":
f = 3s
f = 3*12 = 36 yrs (the father's age)
==============
Cheers,
Stan H.
==============